I am trying to simplify 2^(2/3)+3(2^(2/3) Can you explain how I would do this?
Answer by stanbon(57405)

simplify 2^(2/3)+3(2^(2/3))
Factor out the 2^(2/3) factor:
=2^(2/3) [1 + 3]
=2^(2/3)[2^2]
=2^((2/3)+2))
=2^(8/3)
=============
Cheers,
San H.